OpenStudy (anonymous):

The distances in space are measured in light-years. The distance from Earth to a star is 5.9 × 1013 kilometers. What is the distance, in light-years, from Earth to the star (1 light-year = 9.46 × 1012 kilometers)?

OpenStudy (anonymous):

please help will give medal and fan

OpenStudy (anonymous):

Well so you want to find the answer to 9.46 x 1012 and 5.9 x 1013

OpenStudy (mrnood):

you need to divide the distance to the star by th edistance of a light year: \[\frac{ 5.9 \times 10^{13} }{9.46 \times 10^{12} }\]

OpenStudy (mrnood):

@jackmullen10 those are exponents, not multiplications

OpenStudy (anonymous):

My bad

OpenStudy (mrnood):

@olivia so - are you going to do that division?

OpenStudy (mrnood):

it can be written as \[\frac{ 5.9 }{ 9.46 } \times \frac{ 10^{13} }{ 10^{12}}\]

OpenStudy (mrnood):

so please leave this for the asker to finish.......

OpenStudy (anonymous):

yes sorry I was on the phone with a teacher im solving it now @MrNood